# Heath–Jarrow–Morton meet lifted Heston in energy markets for joint historical and implied calibration

### v1

@@ 2.1.1 Stripping optimization to construct rolling futures contracts
Recall that, except for the first day-ahead contract, a rolling futures contract with time to delivery $T_{s}$ and delivery duration $T_{e}-T_{s}$ is typically not exchanged in standard markets [^5] Although at some specific observation dates $t_{0}$, a rolling future with time to delivery $\delta>0$ and delivery duration $\Delta_{T}>0$ may coincide with the absolute futures contract delivering on $[T_{s},T_{e}]$ if $\begin{cases}\delta=T_{s}-t_{0}\\ \Delta_{T}=T_{e}-T_{s}\end{cases}.$ (2.1) , and therefore its quote cannot be observed directly. Yet, it is possible to construct the quotes of such $P_{\mathrm{hist}}\in\mathbb{N}^{*}$ rolling futures contracts from the ones of the absolute futures contracts observed on the market via a stripped forward curve defined as follows.
@@
$\mathscr{T}_{n}(t_{0})$ :=$:=\left\{t_{0}=T_{0}<T_{1}<\ldots<T_{n}=\bar{T}\right\},$ (2.5)
$\mathscr{T}_{n^{j}}^{j}$ :=$:=\left\{T_{s}^{j}=T_{0}^{j}<T_{1}^{j}<\ldots<T_{n^{j}}^{j}=T_{e}^{j}\right\},\quad j\in\{1,\ldots,P^{\prime}\},$ (2.6)
where $n\in\mathbb{N}^{*}$ (resp. $n^{j}\in\mathbb{N}^{*},\;j\in\{1,\ldots,P^{\prime}\}$) denotes the number of days in the time interval $[t_{0},\bar{T}]$ (resp. in $[T_{s}^{j},T_{e}^{j}],\;j\in\{1,\ldots,P^{\prime}\}$).
@@ 2.2.1 Typical options quoted in power markets
For our numerical experiments, we use the following ten implied volatility smiles observed on the on the $1^{st}$ July 2024 from the German power option market corresponding to eight different underlying futures contracts: Aug 24, Sep 24, Q4 24, Q2 25, Q3 25, Cal 25 Sep, Cal 25 Dec, Cal 26 Mar, Cal 26 Dec, Cal 27 Dec.
2.2.2 Variance swaps deduced from option prices
@@
The calibration will be performed on call and put options written on the futures contracts $F^{i}:=F_{.}(T_{s}^{i},\,T_{e}^{i})$ for $i\in\{1,\ldots,P_{\mathrm{imp}}\}$, where the $i$-th futures contract is identified with its delivery period $[T_{s}^{i},\,T_{e}^{i}]$. For $i\in\{1,\ldots,P_{\mathrm{imp}}\}$, we denote by $N_{i}$ the number of (sorted) maturities $(T_{j}^{i})_{j\in\{1,...,N_{i}\}}$ i.e. the number of smiles, associated to the underlying $F^{i}$. The implied market data contains a family of European call and put option prices
@@
[…]
@@
[…]
@@
3 Calibration methodology for historical covariances and Variance Swaps (VS)
@@ 3.1 Explicit model covariance and VS volatility term structure
Following the definition of the rolling forward contract (1.13) delivering on $[T_{s},T_{e}]$, we define its log returns over a period of $\tau_{d}$-days under the KV approximation by
@@
$t\mapsto a_{X}^{\tau}(t)$ $1$ $e^{\frac{t}{\tau_{S}}}$ $e^{\frac{t}{\tau_{C}}}$ $-\frac{t}{\tau_{C}}e^{\frac{t}{\tau_{C}}}$
$T\mapsto b_{X}^{\tau}(T)$ $1$ $e^{-\frac{T}{\tau_{S}}}$ $\frac{T}{\tau_{C}}e^{-\frac{T}{\tau_{C}}}$ $e^{-\frac{T}{\tau_{C}}}$
$\beta_{X}^{\tau}(T_{e},T_{s})$ $1$ $\frac{\tau_{S}}{T_{e}-T_{s}}\left(e^{\frac{-T_{s}}{\tau_{S}}}-e^{\frac{-T_{e}}{\tau_{S}}}\right)$ $\frac{T_{s}+\tau_{C}}{T_{e}-T_{s}}e^{\frac{-T_{s}}{\tau_{C}}}-\frac{T_{e}+\tau_{C}}{T_{e}-T_{s}}e^{\frac{-T_{e}}{\tau_{C}}}$ $\frac{\tau_{C}}{T_{e}-T_{s}}\left(e^{\frac{-T_{s}}{\tau_{C}}}-e^{\frac{-T_{e}}{\tau_{C}}}\right)$
@@
$C_{2}^{p}$ $-\Big{(}\left(t_{2}-\tau_{C}^{p}\right)e^{\frac{t_{2}}{\tau_{C}^{p}}}-$ $\left(t_{1}-\tau_{C}^{p}\right)e^{\frac{t_{1}}{\tau_{C}^{p}}}\Big{)}$ $-\frac{\tilde{\tau}^{p,k}}{\tau_{C}^{p}}\Big{(}\left(t_{2}-\tilde{\tau}^{p,k}\right)e^{\frac{t_{2}}{\tilde{\tau}^{p,k}}}-$ $\left(t_{1}-\tilde{\tau}^{p,k}\right)e^{\frac{t_{1}}{\tilde{\tau}^{p,k}}}\Big{)},$ $\tilde{\tau}^{p,k}:=\frac{\tau_{S/C}^{k}\tau_{C}^{p}}{\tau_{S/C}^{k}+\tau_{C}^{p}}$ $\frac{\tilde{\tau}^{p,k}}{\tau_{C^{a}}\tau_{C^{b}}}\Big{(}\left(2\left(\tilde{\tau}^{p,k}\right)^{2}-2\tilde{\tau}^{p,k}t_{2}+t_{2}^{2}\right)e^{\frac{t_{2}}{\tilde{\tau}^{p,k}}}-$ $\left(2\left(\tilde{\tau}^{p,k}\right)^{2}-2\tilde{\tau}^{p,k}t_{1}+t_{1}^{2}\right)e^{\frac{t_{1}}{\tilde{\tau}^{p,k}}}\Big{)}$, $\tilde{\tau}^{p,k}:=\frac{\tau_{C}^{k}\tau_{C}^{p}}{\tau_{C}^{k}+\tau_{C}^{p}}$
@@ 3.3 Solver specification
    Note that the loss functional (3.23) is expressed in terms of the state variables, and for each $j\in\{0,\ldots,N_{c}\}$, the $j^{th}$ $C$-factor has two state variables sharing the same parameter $\sigma_{C,j}$ and the same Brownian motion so that we need to impose the following $2N_{c}$ equality constraints on the $x$ variables (3.14) ordered as ($\left(\left\{L,\left(S_{i}\right)_{i\in\left\{0,\ldots,N_{s}\right\}},\left((C_{1,j},C_{2,j})\right)_{i\in\left\{0,\ldots,N_{c}\right\}}\right\}\right)$ such that
@@
Once an optimizer $\left(\hat{a},\hat{x}(\hat{a})\right)$ of the iterative optimization problem (3.27) has been reached, then we extract the optimal parameters $\hat{\tau}$ injecting $\hat{a}$ into (3.29), and by positive-definiteness of the calibrated $L$-$S$-$C$ factors’ covariance $\left(\hat{x}_{i,j}\right)_{i,j\in I}$ (extracted from an optimal solution $\hat{x}(\hat{a})$ of (B.7), with $I$ denoting the set of indices relative to $\left\{L,S,C_{1}\right\}$ state variables, of cardinal $N$), we get
@@ 4.1 Fourier option pricing for implied calibration
In this section, we describe the vanilla options pricing techniques used in the implied calibration for a futures contract $\widetilde{F}_{.}=\widetilde{F}_{.}(T_{s},T_{e})$ with arbitrary delivery period $[T_{s},\,T_{e}]$ in our model, where $\widetilde{F}$ is given by (1.18) following the KV approximation.
@@ 4.2 Exact calibration correction of the VS volatility term structure
The VS term structure calibration consists in matching model and market log-contract prices respectively given by (1.21) and (2.10), i.e. for any IV smile on a futures contract with delivery period $[T_{s},\,T_{e}]$ and maturity $T$, one should ensure that
@@
  • either adjusting the function $T\mapsto g(T)$ on the interval $[T_{s},\,T_{e}]$;
@@
For these reasons, the function $t\mapsto h(t)$ needs to be introduced: we could fix the value of $h$ on $[0,\,T_{1}]$ to satisfy (4.9) for the smile on Cal 25 with maturity $T_{1}$, then adjust $h$ on $[T_{1},T_{4}]$ to ensure (4.9) holds its smile with maturity $T_{4}$. Alternatively, one could calibrate the value of $h$ on $[0,T_{1}]$ to fit (4.9) for the smile on Q4 24 while the smiles on monthly futures are calibrated with the function $g$. However, two problems arise.
@@
    smiles on Oct 24, Nov 24, Dec 24, and Cal 25 with maturity $T_{1}$ are calibrated with the function $g$,
@@
    smiles on Q4 24 and on Cal 25 with maturity $T_{4}$ are calibrated with the function $h$.
Second, once the monthly contracts are calibrated with the function $g$, any modification of the function $h$ (up to the smile maturity) made during the calibration of smile on Q4 24, will impact the log-contract price corresponding to the monthly futures, and vice-versa. This is simply due to the fact that the log-contract price in (4.9) depends both on $g$ on $[T_{s},\,T_{e}]$ and on $h$ on $[0,T]$.
@@
We strip the function $h$ to match the variance swap prices implied from the volatility smiles in (2.9) and the variance swap prices given by the model (1.21) for futures contracts $F^{i}$, for each $i\in\{1,\ldots,P_{\mathrm{imp}}\}$. Fixing $j\in\{1,\ldots,N_{i}\}$ referring to the $j$-th smile of the $i$-th contract, we set $h$ on $[T_{j,\mathrm{prev}}^{i},\,T_{j}^{i})$ equal to
${h(t)}:=\sqrt{\frac{\mathrm{VS}_{T_{j}^{i}}^{i}-\mathrm{VS}_{T_{j,\mathrm{prev}}}^{i}}{\displaystyle\int_{T_{j,\mathrm{prev}}}^{T_{j}^{i}}\Sigma^{\top}_{s}(T_{s}^{i},T_{e}^{i})R\Sigma_{s}(T_{s}^{i},T_{e}^{i}){\rm d}s}},\quad t\in[T_{j,\mathrm{prev}},\,T_{j}^{i}),$ (4.10)
where $T_{j,\mathrm{prev}}$ is the previous smile maturity in the sorted list of all maturities and
$\mathrm{VS}_{T_{j,\mathrm{prev}}}^{i}:=\int_{0}^{T_{j,\mathrm{prev}}}h^{2}(s)\Sigma^{\top}_{s}(T_{s}^{i},T_{e}^{i})R\Sigma_{s}(T_{s}^{i},T_{e}^{i})\,{\rm d}s$
@@
The discontinuity points of $g$ coincide with contracts’ delivery start and end dates $(T_{s}^{i})_{i\in 1,...,P_{\mathrm{imp}}},\;(T_{e}^{i})_{i\in 1,...,P_{\mathrm{imp}}}$. We adopt the following notation for the values of $g$: if the function $g$ is constant on a delivery period $[T_{s}^{i},T_{e}^{i}]$, we will denote its value by $\bar{g}_{i}$. Otherwise, if other delivery periods of futures contracts are included in $[T_{s}^{i},T_{e}^{i}]$, we will denote by $\bar{g}_{i-}$ the value of $g$ on $[T_{s}^{i},T_{e}^{i}]\setminus\cup_{k\in\mathscr{J}_{i}}[T_{s}^{k},T_{e}^{k}]$, where $\mathscr{J}_{i}$ stands for the indices of futures contracts which smiles are calibrated with the function $g$ and which delivery periods are included in $[T_{s}^{i},T_{e}^{i}]$. We also denote by $\hat{\Sigma}_{.}$ the volatility of a futures contract delivering over a set $I$ with $g\equiv 1$ such that
@@
For the contracts $F^{i}$ not calibrated yet, i.e. such that $\mathscr{J}_{i}\not=\varnothing$, the function $g$ should be chosen
@@
with the deterministic volatility component given by (4.12) with $\{\bar{g}_{k}\}_{k\in\mathscr{J}_{i}}$ already calibrated by (4.14). Note that the set $[T_{s}^{i},T_{e}^{i}]\setminus\bigcup_{k\in\mathscr{J}_{i}}[T_{s}^{k},T_{e}^{k}]$ is non-empty thanks to the linear independence condition imposed on the smiles division.
@@
while $||{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g^{n}}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g^{n-1}}||_{\infty}>\epsilon$ do
@@
Denoting one iteration of this algorithm by $\psi:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g^{n}}\mapsto{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g^{n+1}}$, we note that model is calibrated if and only if ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g}$ is a fixed-point of this mapping, i.e. $\psi({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g}$.
@@
[…]

### v2

@@ 2.1.1 Stripping optimization to construct rolling futures contracts
Recall that, except for the first day-ahead contract, a rolling futures contract with time to delivery $T_{s}$ and delivery duration $T_{e}-T_{s}$ is typically not exchanged in standard markets [^5] Although at some specific observation dates $t_{0}$, a rolling future with time to delivery $\delta>0$ and delivery duration $\Delta_{T}>0$ may coincide with the absolute futures contract delivering on $T_{s},T_{e}]$ if $\begin{cases}\delta=T_{s}-t_{0}\\ \Delta_{T}=T_{e}-T_{s}\end{cases}.$ (2.1) , and therefore its quote cannot be observed directly. Yet, it is possible to construct the quotes of such $P_{\mathrm{hist}}\in\mathbb{N}^{*}$ rolling futures contracts from the ones of the absolute futures contracts observed on the market via a stripped forward curve defined as follows.
@@
$\mathscr{T}_{n}(t_{0})$ $:=\left\{t_{0}=T_{0}<T_{1}<\ldots<T_{n}=\bar{T}\right\},$ (2.5)
$\mathscr{T}_{n^{j}}^{j}$ $:=\left\{T_{s}^{j}=T_{0}^{j}<T_{1}^{j}<\ldots<T_{n^{j}}^{j}=T_{e}^{j}\right\},\quad j\in\{1,\ldots,P^{\prime}\},$ (2.6)
where $n\in\mathbb{N}^{*}$ (resp. $n^{j}\in\mathbb{N}^{*},\;j\in\{1,\ldots,P^{\prime}\}$) denotes the number of days in the time interval $t_{0},\bar{T}]$ (resp. in $T_{s}^{j},T_{e}^{j}],\;j\in\{1,\ldots,P^{\prime}\}$).
@@ 2.2.1 Typical options quoted in power markets
For our numerical experiments, we use the following ten implied volatility smiles observed on the on the $1^{st}$ July 2024 from the German power option market corresponding to eight different underlying futures contracts: Aug 24, Sep 24, Q4 24, Q2 25, Q3 25, Cal 25 Sep, Cal 25 Dec, Cal 26 Mar, Cal 26 Dec, Cal 27 Dec. Note that in energy markets, a large part of the volume is traded over the counter (OTC), unlike in equity markets, where options are typically traded via exchanges in a centralized manner. Consequently, there exist many possible sources of option price data. For our study, we use implied volatility data provided by the broker ICAP [^9] https://icap.com/energy-commodities.
2.2.2 Variance swap volatilities deduced from option prices
@@
The calibration will be performed on call and put options written on the futures contracts $F^{i}:=F_{.}(T_{s}^{i},\,T_{e}^{i})$ for $i\in\{1,\ldots,P_{\mathrm{imp}}\}$, where the $i$-th futures contract is identified with its delivery period $T_{s}^{i},\,T_{e}^{i}]$. For $i\in\{1,\ldots,P_{\mathrm{imp}}\}$, we denote by $N_{i}$ the number of (sorted) maturities $(T_{j}^{i})_{j\in\{1,...,N_{i}\}}$ i.e. the number of smiles, associated to the underlying $F^{i}$. The implied market data contains a family of European call and put option prices
@@
  • $\rho_{i}\in[-1,\,1]$ is a skew parameter interpreted as a spot-vol correlation for the contract $F^{i}$;
@@
Due to the presence of multiple surfaces to be parametrized together, we specify individual correlation parameters $\rho_{i}\in[-1,\,1]$ for each contract $F^{i}$, and a common function $\phi$ since we assume the same nature of volatility across all the underlying futures contracts. This choice appears consistent with observed market data and allows us to simultaneously parametrize all the volatility surfaces using a small number of parameters. We set
$\phi(\theta):=\dfrac{\eta}{\theta^{\gamma}(1+\theta)^{1-\gamma}},\quad\gamma\in(0,\,0.5],$
@@
3 Calibration methodology for historical covariances and variance swap volatilities
@@ 3.1 Explicit model covariance and VS volatility term structure
Following the definition of the rolling forward contract (1.13) delivering on $T_{s},T_{e}]$, we define its log returns over a period of $\tau_{d}$-days under the KV approximation by
@@
$t\mapsto a_{X}^{\tau}(t)$ 11 $e^{\frac{t}{\tau_{S}}}$ $e^{\frac{t}{\tau_{C}}}$ $\frac{t}{\tau_{C}}e^{\frac{t}{\tau_{C}}}$
$T\mapsto b_{X}^{\tau}(T)$ 11 $e^{-\frac{T}{\tau_{S}}}$ $\frac{T}{\tau_{C}}e^{-\frac{T}{\tau_{C}}}$ $e^{-\frac{T}{\tau_{C}}}$
$\beta_{X}^{\tau}(T_{e},T_{s})$ 11 $\frac{\tau_{S}}{T_{e}-T_{s}}\left(e^{\frac{-T_{s}}{\tau_{S}}}-e^{\frac{-T_{e}}{\tau_{S}}}\right)$ $\frac{T_{s}+\tau_{C}}{T_{e}-T_{s}}e^{\frac{-T_{s}}{\tau_{C}}}-\frac{T_{e}+\tau_{C}}{T_{e}-T_{s}}e^{\frac{-T_{e}}{\tau_{C}}}$ $\frac{\tau_{C}}{T_{e}-T_{s}}\left(e^{\frac{-T_{s}}{\tau_{C}}}-e^{\frac{-T_{e}}{\tau_{C}}}\right)$
@@
$C_{2}^{p}$ $\Big(\left(t_{2}-\tau_{C}^{p}\right)e^{\frac{t_{2}}{\tau_{C}^{p}}}-$ $\left(t_{1}-\tau_{C}^{p}\right)e^{\frac{t_{1}}{\tau_{C}^{p}}}\Big)$ $\frac{\tilde{\tau}^{p,k}}{\tau_{C}^{p}}\Big(\left(t_{2}-\tilde{\tau}^{p,k}\right)e^{\frac{t_{2}}{\tilde{\tau}^{p,k}}}-$ $\left(t_{1}-\tilde{\tau}^{p,k}\right)e^{\frac{t_{1}}{\tilde{\tau}^{p,k}}}\Big),$ $\tilde{\tau}^{p,k}:=\frac{\tau_{S/C}^{k}\tau_{C}^{p}}{\tau_{S/C}^{k}+\tau_{C}^{p}}$ $\frac{\tilde{\tau}^{p,k}}{\tau_{C^{a}}\tau_{C^{b}}}\Big(\left(2\left(\tilde{\tau}^{p,k}\right)^{2}-2\tilde{\tau}^{p,k}t_{2}+t_{2}^{2}\right)e^{\frac{t_{2}}{\tilde{\tau}^{p,k}}}$- $\left(2\left(\tilde{\tau}^{p,k}\right)^{2}-2\tilde{\tau}^{p,k}t_{1}+t_{1}^{2}\right)e^{\frac{t_{1}}{\tilde{\tau}^{p,k}}}\Big)$, $\tilde{\tau}^{p,k}:=\frac{\tau_{C}^{k}\tau_{C}^{p}}{\tau_{C}^{k}+\tau_{C}^{p}}$
@@ 3.3 Solver specification
    Note that the loss functional (3.23) is expressed in terms of the state variables, and for each $j\in\{0,\ldots,N_{c}\}$, the $j^{th}$ $C$-factor has two state variables sharing the same parameter $\sigma_{C,j}$ and the same Brownian motion so that we need to impose the following $2N_{c}$ equality constraints on the $x$ variables (3.14) ordered as $\left(\left\{L,\left(S_{i}\right)_{i\in\left\{0,\ldots,N_{s}\right\}},\left((C_{1,j},C_{2,j})\right)_{i\in\left\{0,\ldots,N_{c}\right\}}\right\}\right)$ such that
@@
[…]
@@ 4.1 Fourier option pricing for implied calibration
In this section, we describe the vanilla options pricing techniques used in the implied calibration for a futures contract $\widetilde{F}_{.}=\widetilde{F}_{.}(T_{s},T_{e})$ with arbitrary delivery period $T_{s},\,T_{e}]$ in our model, where $\widetilde{F}$ is given by (1.20) following the KV approximation.
@@ 4.2 Exact calibration correction of the VS volatility term structure
The VS term structure calibration consists in matching model and market log-contract prices respectively given by (1.23) and (2.10), i.e. for any IV smile on a futures contract with delivery period $T_{s},\,T_{e}]$ and maturity $T$, one should ensure that
@@
  • either adjusting the function $T\mapsto g(T)$ on the interval $T_{s},\,T_{e}]$;
@@
For these reasons, the function $t\mapsto h(t)$ needs to be introduced: we could fix the value of $h$ on $0,\,T_{1}]$ to satisfy (4.9) for the smile on Cal 25 with maturity $T_{1}$, then adjust $h$ on $T_{1},T_{4}]$ to ensure (4.9) holds its smile with maturity $T_{4}$. Alternatively, one could calibrate the value of $h$ on $0,T_{1}]$ to fit (4.9) for the smile on Q4 24 while the smiles on monthly futures are calibrated with the function $g$. However, two problems arise.
@@
    the function $g$ is calibrated to the smiles on Oct 24, Nov 24, Dec 24, and Cal 25 with maturity $T_{1}$,
@@
    the function $h$ is calibrated to the smiles on Q4 24 and on Cal 25 with maturity $T_{4}$.
Second, once the monthly contracts are calibrated with the function $g$, any modification of the function $h$ (up to the smile maturity) made during the calibration of smile on Q4 24, will impact the log-contract price corresponding to the monthly futures, and vice-versa. This is simply due to the fact that the log-contract price in (4.9) depends both on $g$ on $T_{s},\,T_{e}]$ and on $h$ on $[0,T]$.
@@
We strip the function $h$ to match the variance swap prices implied from the volatility smiles in (2.9) and the variance swap prices given by the model (1.23) for futures contracts $F^{i}$, for each $i\in\{1,\ldots,P_{\mathrm{imp}}\}$. Fixing $j\in\{1,\ldots,N_{i}\}$ referring to the $j$-th smile of the $i$-th contract, we set $h$ on $T_{j,\mathrm{prev}}^{i},\,T_{j}^{i})$ equal to
${h(t)}:=\sqrt{\frac{\mathrm{VS}_{T_{j}^{i}}^{i}-\mathrm{VS}_{T_{j,\mathrm{prev}}^{i}}^{i}}{\displaystyle\int_{T_{j,\mathrm{prev}}^{i}}^{T_{j}^{i}}\Sigma^{\top}_{s}(T_{s}^{i},T_{e}^{i})R\Sigma_{s}(T_{s}^{i},T_{e}^{i}){\rm d}s}},\quad t\in[T_{j,\mathrm{prev}}^{i},\,T_{j}^{i}),$ (4.10)
where $T_{j,\mathrm{prev}}^{i}$ is the previous smile maturity in the sorted list of all maturities and
$\mathrm{VS}_{T_{j,\mathrm{prev}}^{i}}^{i}:=\int_{0}^{T_{j,\mathrm{prev}}^{i}}h^{2}(s)\Sigma^{\top}_{s}(T_{s}^{i},T_{e}^{i})R\Sigma_{s}(T_{s}^{i},T_{e}^{i})\,{\rm d}s$
@@
The discontinuity points of $g$ coincide with contracts’ delivery start and end dates $(T_{s}^{i})_{i\in 1,...,P_{\mathrm{imp}}},\;(T_{e}^{i})_{i\in 1,...,P_{\mathrm{imp}}}$. We adopt the following notation for the values of $g$: if the function $g$ is constant on a delivery period $T_{s}^{i},T_{e}^{i}]$, we will denote its value by $\bar{g}_{i}$. Otherwise, if other delivery periods of futures contracts are included in $T_{s}^{i},T_{e}^{i}]$, we will denote by $\bar{g}_{i-}$ the value of $g$ on $T_{s}^{i},T_{e}^{i}]\setminus\cup_{k\in\mathscr{J}_{i}}[T_{s}^{k},T_{e}^{k}]$, where $\mathscr{J}_{i}$ stands for the indices of futures contracts which smiles are calibrated with the function $g$ and which delivery periods are included in $T_{s}^{i},T_{e}^{i}]$. We also denote by $\hat{\Sigma}_{.}$ the volatility of a futures contract delivering over a set $I$ with $g\equiv 1$ such that
@@
For the contracts $F^{i}$ not calibrated yet, i.e. such that $\mathscr{J}_{i}\not$$\varnothing$, the function $g$ should be chosen
@@
with the deterministic volatility component given by (4.12) with $\{\bar{g}_{k}\}_{k\in\mathscr{J}_{i}}$ already calibrated by (4.14). Note that the set $T_{s}^{i},T_{e}^{i}]\setminus\bigcup_{k\in\mathscr{J}_{i}}[T_{s}^{k},T_{e}^{k}]$ is non-empty thanks to the linear independence condition imposed on the smiles division.
@@
while ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g^{n}}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g^{n-1}}||_{\infty}>\epsilon$ do
@@
[…]
@@
$\left(\dfrac{1}{4}\right)^{2}\int_{0}^{T_{\mathrm{Dec}}}h(s)^{2}\Sigma_{s}^{\top}(\mathrm{Q2\;25})R\Sigma_{s}(\mathrm{Q2\;25})\,{\rm d}s<\mathrm{VS}^{\mathrm{Cal\;25}}_{T_{\mathrm{Dec}}}.$ (4.20)